OpenStudy (anonymous):

A rectangular garden has a width of 5 feet and a length of 10 feet. If an equal amount is added to both the width and the length, the area is increased to 84 square feet. What is this amount?

OpenStudy (anonymous):

(5+x)(10+x)=84

OpenStudy (anonymous):

x^2+15x+50-84=0

OpenStudy (anonymous):

x^2+15x-34=0

OpenStudy (anonymous):

(x+17)(x-2)=0

OpenStudy (anonymous):

the amount is 2 we reject -17 cause it doesnt make sense

OpenStudy (anonymous):

TADA!

OpenStudy (anonymous):

WOW, thanks!

OpenStudy (anonymous):

np